Calculate the final concentration of the solution when water is added to prepare each of the following solutions.

65.0 mL of a 22.0 M HCl solution is diluted to 520. mL

50.0 mL of a 21.0 %(m/v) NH4Cl solution is diluted to 175 mL

5.00 mL of a 9.50 M KOH solution is diluted to 90.0 mL
Answer #1

1)

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 22 M

V1 = 65 mL

V2 = 520 mL

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (22*65)/520

M2 = 2.75 M

Answer: 2.75 M

2)

Given:

M1 = 21 %

V1 = 50 mL

V2 = 175 mL

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (21*50)/175

M2 = 6.00 %

Answer: 6.00 %

3)

Given:

M1 = 9.5 M

V1 = 5 mL

V2 = 90 mL

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (9.5*5)/90

M2 = 0.5278 M

Answer: 0.528 M